# Small-scale magnetohydrodynamic dynamos: from deterministic chaos to turbulence

### v1

@@ IV.1 Dissipative MHD invariants
Analogously, using another form of the Kármán-Howarth equation suggested in the paper Ref. wan , it can be shown that the magnetic analogy of the Birkhoff-Saffman integral is also a dissipative (adiabatic) invariant
@@ VI Geodynamo
The time-averaged spatial spectral data at the spheres of different radii $R$ were provided in the paper Ref. sjna for the case with the most vigorous motion: $Re=5140$, $Ra=2.54\times 10^{12}$, $Ro=5.1\times 10^{-4}$, $E=10^{-7}$, $Pr=1$, $Re_{m}=514$, and $Pr_{m}=0.1$. The system is helical, and for this case the bulk relative helicity and helicity at the boundaries was also rather strong.
It should be noted that the spectral data were provided in the Ref. sjna as a dependence of the energy density spectra $E(l)$ on the spherical harmonic degree $l$. On a sphere of radius $R$ the wavenumber $k$ is related to $l$ as
$kR=\sqrt{l(l+1)}$ (41)41( 41 )
Figure 11 shows geomagnetic power spectrum calculated using local subgrids from World Magnetic Anomaly Map of the National Geophysics Data Center (WMAM) maus1 ). The data were detrended and azimuthally averaged, and the mean field was subtracted by author of the Ref. maus2 (the spectral data were taken from Fig. 4 of the Ref. maus2 ). The land and ocean averages were considered separately (here $R$ is the Earth’s radius).
The dashed curve indicates the helical spectral law Eq. (9) for the ocean data, whereas at the land the spectral law Eq. (21) indicates the magneto-inertial range of scales.
In the paper Ref. maus1 a large-scale globally averaged total intensity spectrum computed using the spherical harmonic coefficients of the magnetic potential from a C89 global geomagnetic field model C89 cain was also presented. It is interesting to compare the spectral data with the above suggested spectra. Figure 12 shows the spectral data taken from Fig. 2 of the Ref. maus1 . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
The C89 model used the data available at the time of its creation. Therefore it is useful to look at the modern geomagnetic data including, in particular, the low Earth orbit space-based observations. An example of such global geomagnetic energy spectrum is given in Fig. 13 (in the semi-log scales). The spectral data were taken from Fig. 13 (the main field) of a recent paper Ref. styp . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
@@ VII Solar small-scale MHD dynamos
Both global (large-scale) and local (small-scale) solar dynamos have been vigorously studied in the last decades. Chaotic nonlinear global dynamo mechanisms attract special attention (see, for instance, Refs. cha , ber2 and references therein). Chaotic small-scale dynamo mechanisms (localized mainly in the near-surface solar layer) attract less attention. It is a little strange since the small-scale solar dynamos are usually related in the literature to the quiet Sun magnetic fields (see, for instance, Ref. remp and references therein). As it was already mentioned above the large- and small-scale dynamos can well coexist (see, for instance, Fig. 3) and the highly active solar regions can exhibit chaotic behavior ber3 .
It also can be seen in the Fig. 2 showing the time evolution of the magnetic energy spectra in a large emerging solar active region AR NOAA 11726. The spectral data were taken from Fig. 2a of a recent paper Ref. kka . The measurements were provided by Helioseismic and Magnetic Imager located on board of the Solar Dynamics Observatory (SDO/HMI). These spectra are well-fitted by the sequence of the spectral laws (the dashed curves): starting from the deterministic chaos Eq. (2) ($\beta=1$), then the dissipative distributed chaos Eq (31) ($\beta=3/4$), then the helical distributed chaos Eq. (9) ($\beta=1/2$), and then the magneto-inertial range of scales Eq. (21) ($\beta=1/4$, a precursor of the hard turbulence).
@@ VII.1 Numerical simulations
Figure 14: Magnetic energy spectra for a photosphere layer at different times of the dynamo evolution.
Figure 15: Magnetic energy spectra averaged over a vertical range between $z=0$ km and $z=-200$ km, and over time. The dissipative and non-dissipative cases correspond to the bottom and top curves respectively. The spectra are vertically shifted for clarity.
Figure 16: Magnetic energy spectra in undisturbed photosphere: CH - coronal hole, QS - quiet Sun region, SG - supergranular network, MF - magnetic field.
@@ VII.2 Solar observations
Figure 16 shows magnetic energy spectra in an undisturbed (outside active regions) photosphere’s area computed using the results of measurements produced with the Helioseismic and Magnetic Imager onboard Solar Dynamic Observatory on June 19, 2017, in the center of the solar disk. The spectral data were taken from Fig. 4 of a paper Ref. ak . The solar area consists of three subregions: a coronal hole (CH), a quiet Sun region (QS), and a supergranular network (SG). On the whole, it was an area of weak magnetic fields, but the CH, QS, and SG subregions had comparatively different intensities of their magnetic fields: CH - weakest, QS - moderate, and SG - strongest. The super-granular network (SG subregion) was an area of decayed several rotations ago NOAA 12242–12259 active regions.
The dashed curves indicate the best fit of the spectra with the exponent Eq. (2) ($\beta=1$, deterministic chaos in the CH magnetic field), with the Eq. (33) ($\beta=5/6$, dissipative distributed chaos dominated by the magnetic Loitsianskii invariant in the QS moderate magnetic filed), and with the Eq. (31) ($\beta=3/4$, dissipative distributed chaos dominated by the magnetic Birkhoff-Saffman invariant in the SG strongest magnetic field).
@@
To take into account a mean magnetic field (if it is present) one should replace the energy dissipation rate $\varepsilon$ in the Eq. (16) by $(\varepsilon\widetilde{B}_{0})$ (where $\widetilde{B}_{0}=B_{0}/\sqrt{\mu_{0}\rho}$ is normalized mean magnetic field - in the Alfvenic units, with the same dimension as velocity) ir .
Then dimensional considerations result in
@@
i.e. the parameter $\alpha=1/8$, and from Eq. (20) we obtain $\beta=1/5$
@@
Figure 17 shows the magnetic energy spectrum for a fully developed active region NOAA 09077. The spectral data were taken from Fig. 5 of a paper Ref. hew . The solar magnetogram images used for the calculation of the spectra were obtained with the Michelson Doppler Imager (MDI) onboard the Solar and Heliospheric Observatory (SOHO) spacecraft. The dashed curve in Fig. 17 indicates the best fit of the spectrum with the Eq. (43) ($\beta=1/5$), corresponding to the magneto-inertial range under the strong influence of a mean magnetic field.
Figure 18 shows the magnetic energy spectra for a rapidly emerging active region NOAA 10488 at different times of its evolution. The spectral data were taken from Fig. 8 of a paper Ref. hew . The solar magnetogram images used for the calculation of the spectra were also obtained with the MDI. An M-class flare and two C-class flares occurred in the active region, and also an active region NOAA 10493 appeared near the AR NOAA 10488 during the period of 27-28 October 2003.
Figure 17: Magnetic energy spectrum for a fully developed active region NOAA 09077.
Figure 18: Magnetic energy spectra for a rapidly emerging solar active region AR NOAA 10488 at different times of its evolution. The spectra are vertically shifted for clarity.
The dashed curves in Fig. 17 indicate the best fit of the spectra with the Eq. (33) ($\beta=5/6$) and Eq. (31) ($\beta=3/4$), corresponding to the dissipative distributed chaos dominated by the magnetic Loitsianskii and Birkhoff-Saffman invariants at the earlier stages of the AR NOAA 10488 emergence; and with the Eq. (21) ($\beta=1/4$) and Eq. (43) ($\beta=1/5$), corresponding to the magneto-inertial range of scales at the mature stages of the AR NOAA 10488 emergence (cf Fig. 1).
@@ VIII Conclusions
In the frames of this approach the results of numerical simulations are in quantitative agreement with the geophysical and solar observations despite the considerable differences in the scales and physical parameters.

### v2

@@ IV.1 Dissipative MHD invariants
Analogously, using another form of the Kármán-Howarth equation suggested in the paper Ref. wan , it can be shown that the magnetic analogy of the Birkhoff-Saffman integral is also a dissipative (adiabatic) invariant ber4
@@ VI Geodynamo
The time-averaged spatial spectral data at the spheres of different radii $R$ were provided in the paper Ref. sjna for the case with the most vigorous motion: $Re=5140$, $Ra=2.54\times 10^{12}$, $Ro=5.1\times 10^{-4}$, $E=10^{-7}$, $Pr=1$, $Re_{m}=514$, and $Pr_{m}=0.1$. The system is helical, and for this case the bulk relative helicity and helicity at the boundaries were also rather strong.
It should be noted that the spectral data were provided in Ref. sjna as a dependence of the energy density spectra $E(l)$ on the spherical harmonic degree $l$. On a sphere of radius $R$ the wavenumber $k$ is related to $l$ as
$kR=\sqrt{l(l+1)}$ (41)41( 41 )
Figure 11 shows the geomagnetic power spectrum calculated using local subgrids from the World Magnetic Anomaly Map of the National Geophysics Data Center (WMAM) maus1 ). The data were detrended and azimuthally averaged, and the mean field was subtracted by the author of the Ref. maus2 (the spectral data were taken from Fig. 4 of the Ref. maus2 ). The land and ocean averages were considered separately (here $R$ is the Earth’s radius). The dashed curve indicates the helical spectral law Eq. (9) for the ocean data, whereas at the land the spectral law Eq. (21) indicates the magneto-inertial range of scales.
@@
In the paper Ref. maus1 a large-scale globally averaged total intensity spectrum computed using the spherical harmonic coefficients of the magnetic potential from a C89 global geomagnetic field model C89 cain was also presented. It is interesting to compare the spectral data with the above-suggested spectra. Figure 12 shows the spectral data taken from Fig. 2 of the Ref. maus1 . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
The C89 model used the observational data available at the time of its creation. Therefore it is useful to look at the modern geomagnetic data including, in particular, the low Earth orbit space-based observations. An example of such a global geomagnetic energy spectrum is given in Fig. 13 (in the semi-log scales). The spectral data were taken from Fig. 13 (the main field) of a recent paper Ref. styp . The dashed curves indicate the helical spectrum Eq. (9) (here $R$ is the Earth’s radius).
@@ VII Solar small-scale MHD dynamos
Both global (large-scale) and local (small-scale) chaotic/turbulent solar dynamos have been vigorously studied in the last decades (see, for instance, Refs. cha , ber2 ,remp and references therein). It is believed that the chaotic/turbulent small-scale dynamo mechanisms are localized mainly in the near-surface solar layer. As it was already mentioned above the large- and small-scale dynamos can well coexist (see, for instance, Fig. 3) and both the quiet Sun and the highly active solar regions remp ber3 can exhibit chaotic behavior.
@@ VII.2 Solar observations
Figure 14: Magnetic energy spectra for a photospheric layer at different times of the dynamo evolution (DNS).
Figure 15: Magnetic energy spectra averaged over a vertical range between $z=0$ km and $z=-200$ km, and over time. The dissipative and non-dissipative cases correspond to the bottom and top curves respectively (DNS). The spectra are vertically shifted for clarity.
Figure 16: Magnetic energy spectra for synoptic maps of the early rising (bottom) and early declining (top) phases of solar cycle 24.
Figure 17: Magnetic energy spectra in undisturbed photosphere: CH - coronal hole (bottom), QS - quiet Sun region, SG - supergranular network (top), MF - magnetic field.
Figure 16 shows the magnetic energy spectra for synoptic maps of the early rising and early declining phases of solar cycle 24. In the time of the early declining phase both magnetic energy and magnetic helicity have their maxima, whereas during the early rising phase, the magnetic energy has its minimum singh . The spectral data were taken from Figs. 4.2b and 4.3b of Ref. prabhu . The data used for computing the spectrum of the early declining phase were obtained by the Helioseismic and Magnetic Imager onboard Solar dynamics observatory, whereas the data used for computing the spectrum of the early rising phase were obtained by the Vector Spectromagnetograph on the ground-based Synoptic Optical Long-term Investigations of the Sun telescope.
The dashed curves indicate the best fit of the spectra with Eq. (31) ($\beta=3/4$, dissipative distributed chaos dominated by the magnetic Birkhoff-Saffman invariant) for the early rising phase and Eq. (21) ($\beta=1/4$, the magneto-inertial range of scales) for the early declining phase when both magnetic energy and magnetic helicity have their maxima.
Figure 17 shows magnetic energy spectra in an undisturbed (outside active regions) photospheric area computed using the results of measurements produced with the Helioseismic and Magnetic Imager onboard Solar Dynamic Observatory on June 19, 2017, in the center of the solar disk. The spectral data were taken from Fig. 4 of a paper Ref. ak . The solar area consists of three subregions: a coronal hole (CH), a quiet Sun region (QS), and a supergranular network (SG). On the whole, it was an area of weak magnetic fields, but the CH, QS, and SG subregions had comparatively different intensities of their magnetic fields: CH - weakest, QS - moderate, and SG - strongest. The super-granular network (SG subregion) was an area of decayed several rotations ago NOAA 12242–12259 active regions.
The dashed curves indicate the best fit of the spectra with the exponent Eq. (2) ($\beta=1$, deterministic chaos in the CH weak magnetic field), with the Eq. (33) ($\beta=5/6$, dissipative distributed chaos dominated by the magnetic Loitsianskii invariant in the QS moderate magnetic filed), and with the Eq. (31) ($\beta=3/4$, dissipative distributed chaos dominated by the magnetic Birkhoff-Saffman invariant in the SG strongest magnetic field).
@@
Figure 2 shows the evolution of the magnetic energy spectra for a large emerging solar active region NOAA 11726. The spectral data were taken from Fig. 2a of the recent paper Ref. kka . The measurements were provided by Helioseismic and Magnetic Imager located on board of the Solar Dynamics Observatory. These spectra are well-fitted by the sequence of the spectral laws (the dashed curves): starting from the deterministic chaos Eq. (2) ($\beta=1$), then the dissipative distributed chaos Eq (31) ($\beta=3/4$), then the helical distributed chaos Eq. (9) ($\beta=1/2$), and then the magneto-inertial range of scales Eq. (21) ($\beta=1/4$, a precursor of the hard turbulence.
Figure 18 shows the evolution of the magnetic energy spectra for a smaller emerging solar active region NOAA 11781. The spectral data were taken from Fig. 2b of the same paper Ref. kka . The measurements were also provided by Helioseismic and Magnetic Imager located on board the Solar Dynamics Observatory. These spectra are well-fitted by the sequence of the analogous spectral laws (the dashed curves) and only the most energetic magneto-inertial range of scales Eq. (21) ($\beta=1/4$) is absent here (cf Fig. 2).
Figure 18: Evolution of the magnetic energy spectra for an emerging solar active region NOAA 11781.
Figure 19: Evolution of the magnetic energy spectra for an emerging solar active region NOAA 12219.
VII.3 The role of a mean magnetic field
To take into account a mean magnetic field (if it is present) in the case of a magneto-inertial range of scales one should replace the energy dissipation rate $\varepsilon$ in the Eq. (16) by $(\varepsilon\widetilde{B}_{0})$ (where $\widetilde{B}_{0}=B_{0}/\sqrt{\mu_{0}\rho}$ is normalized mean magnetic field - in the Alfvenic units, with the same dimension as velocity) ir , and the $\varepsilon_{h}$ by the modified magnetic helicity (Eq. 4) dissipation rate $\varepsilon_{\hat{h}}$.
The dimensional considerations result in
@@ VII.3 The role of a mean magnetic field
i.e. the parameter $\alpha=1/8$. Then from Eq. (20) we obtain $\beta=1/5$
@@
Figure 19 shows the evolution of the magnetic energy spectra for an emerging solar active region NOAA 12219. The spectral data were taken from Fig. 1 of a recent paper Ref. kak . The solar magnetogram images used for the calculation of the spectra were obtained by the Helioseismic and Magnetic Imager onboard the Solar Dynamics Observatory spacecraft. The time $t_{0}$ corresponds to the onset of the active region emergence, $t_{1}$ corresponds to the time when only the first imprints of the emerging active region can be seen in the magnetogram, the time $t_{2}$ corresponds to the peak of magnetic flux.
These spectra are fitted by the sequence of the spectral laws (the dashed curves): starting from the deterministic chaos Eq. (2) ($\beta=1$), then the dissipative distributed chaos Eq (31) ($\beta=3/4$), and then the magneto-inertial range of scales Eq. (43) ($\beta=1/5$), corresponding to the magneto-inertial range under the strong influence of a mean magnetic field.
Figure 20 shows the evolution of the magnetic energy spectra for a rapidly emerging active region NOAA 10488. The spectral data were taken from Fig. 8 of a paper Ref. hew . The solar magnetogram images used for the calculation of the spectra were obtained by the Michelson Doppler Imager onboard the Solar and Heliospheric Observatory. An M-class flare and two C-class flares occurred in the active region, and also an active region NOAA 10493 appeared near the AR NOAA 10488 during the period of 27-28 October 2003.
The dashed curves in Fig. 20 indicate the best fit of the spectra with the Eq. (33) ($\beta=5/6$) and Eq. (31) ($\beta=3/4$), corresponding to the dissipative distributed chaos dominated by the magnetic Loitsianskii and Birkhoff-Saffman invariants at the earlier stages of the AR NOAA 10488 emergence; and with the Eq. (21) ($\beta=1/4$) and Eq. (43) ($\beta=1/5$), corresponding to the magneto-inertial range of scales at the mature stage of the AR NOAA 10488 emergence (cf Fig. 2).
Figure 20: Evolution of the magnetic energy spectra for a rapidly emerging solar active region AR NOAA 10488. The spectra are vertically shifted for clarity.
@@ VIII Conclusions
In the frames of this approach, the results of numerical simulations are in quantitative agreement with the geophysical and solar observations despite the considerable differences in the scales and physical parameters.